OpenStudy (anonymous):

the roots of the equation f(x)=0 are 1 and -2. The roots of f(2x)=0 are.... A. 1 and -2 B. 1/2 and -1 C. -1/2 and 1 D. 2 and -4 E. -2 and 4

OpenStudy (ybarrap):

Roots mean for what values of x make f(x) zero. They give you that when x =1 or x=-2 that f(x) = 0. This means that f(1)=0 and f(-2)=0. If f(x) = x-1, then if x=1, f(x)=0. Do you see why?

OpenStudy (anonymous):

So would the answer be C? @ybarrap

OpenStudy (ybarrap):

No. Why do you think that?

OpenStudy (ybarrap):

We haven't dealt with f(2x) yet. We only know if f(x)=x-1, then f(1)=0. We still need to add another team to make this same function equal zero at negative 2. That would be if f(x)=(x+2). Now how do we combine these two results so that f(1)=0 and f(2)=0? Once we're done with that, we can set y=2x to find f(y)=0 and immediately get our answer.

OpenStudy (ybarrap):

f(x)=(x-1)(x+2) Does this satisfy f(1)=f(-2)=0?

OpenStudy (anonymous):

Yes!!

OpenStudy (ybarrap):

But we need f(2x). This is easy now. f(2x)=(2x-1)(2x+2)=0. What values of x make this true?

OpenStudy (anonymous):

x=1/2 and x=-1

OpenStudy (ybarrap):

Perfect!

OpenStudy (anonymous):

Thanks!!!

OpenStudy (ybarrap):

yw